# M-LLM Based Video Frame Selection for Efficient Video Understanding

### v1

@@ Abstract
Recent advances in Multi-Modal Large Language Models (M-LLMs) show promising results in video reasoning. Popular Multi-Modal Large Language Model (M-LLM) frameworks usually apply naive uniform sampling to reduce the number of video frames that are fed into an M-LLM, particularly for long context videos. However, it could lose crucial context in certain periods of a video, so that the downstream M-LLM may not have sufficient visual information to answer a question. To attack this pain point, we propose a light-weight M-LLM-based frame selection method that adaptively select frames that are more relevant to users’ queries. In order to train the proposed frame selector, we introduce two supervision signals (i) Spatial signal, where single frame importance score by prompting a M-LLM; (ii) Temporal signal, in which multiple frames selection by prompting Large Language Model (LLM) using the captions of all frame candidates. The selected frames are then digested by a frozen downstream video M-LLM for visual reasoning and question answering. Empirical results show that the proposed M-LLM video frame selector improves the performances various downstream video Large Language Model (video-LLM) across medium (ActivityNet, NExT-QA) and long (EgoSchema, LongVideoBench) context video question answering benchmarks.
@@ 1 Introduction
Figure 1: An example of our video frame selection for video QA. Compared to uniform sampling, our select has higher hit rate.
@@
The rise of Large Language Model (LLM) has revolutionized numerous domains in Artificial Intelligence (AI[29, 3, 36, 39]. In the past year, Multi-Modal Large Language Model (M-LLM) has significantly improved the performances of Vision-Language Models (VLM) to an unprecedented level in tasks such as image captioning and Visual Question Answering (VQA[2, 12, 4, 41]. Video QA task requires a model to understand consecutive images along the time to answer a question, raises new challenges to the current M-LLM. One of most critical challenge is the length of the video context, where a model needs to comprehend all frames in a video.
@@
To balance the trade-off between the capability of understand all video frames and the available context length in a specific M-LLM, conventional practices [48, 54] rely on uniform sampling of frames, wherein frames are extracted at pre-defined intervals regardless of their relevance to the specific questions. Although this method ensures the full coverage of a video in time-axis, it also introduces insufficient visual information, as many frames, which may be irrelevant or redundant are included, on the other some important frames are ignored. This dilutes the model’s ability to focus on key events in the video and sometimes increases the computational cost. Generally in video QA [52, 21], some specific frames are more likely to contain information pertinent to the question, this one-size-fits-all approach limits both the performance and practicality of M-LLMs. The need for adaptive frame selection is becoming important, especially when working with long videos or resource-limited environments.
@@
To address these limitations, a most straightforward idea is to select frames instead of uniform sampling [50, 30]. By focusing on the frames that helps answer the question most, we can significantly reduce the visual context that an M-LLM needs to be processed without sacrificing the quality of video understanding. Illustrated as an example in Figure 1, frames that contains the most informative visuals can help downstream model to answer the question. To implement the above idea, we propose a light weight frame selector that employs fine-tuned version of an LLM. It makes use of M-LLM’s capability of multi-modality understanding to effectively capture the relationship between video frames and the question. We introduce two design choices to make the framework lightweight. First, we demonstrate that small LLMs are capable to understand complicated user questions. Second, we compress per video frame tokens to balance the long context trade-off.
@@
Although video question-answering tasks often require a large number of tokens per frame to capture content details, we hypothesize that determining frame importance does not require excessive tokens. Instead of leveraging all visual tokens of the frames, we apply an aggressive pooling-based token reduction on each video frame, which significantly improves the computational efficiency and increases the number of frames that the LLM selector can ingest.
@@
We also propose a method to train our LLM-based video frame selector. Given the fact that there are very limited resources of well-maintained video frame selection datasets for video QA, we are not able to simply apply supervised training. To overcome this limitation, we use two pseudo-labeling strategies to estimate the importance of each frame in a video. First is the spatial understanding, where we use a video QA question to prompt a well-trained M-LLM, which generates the importance score for each frame accordingly. However, the context length of an M-LLM is limited, it impossible to jointly estimate the importance scores of all frames from the temporal perspective. We then take advantage of an LLM to select top-k relevant frames in a video using their captions. Using captions rather than visual tokens, an LLM is able to understand a larger number of frames at a time. We combine these two approaches to obtain pseudo-labels indicating frame importance in given a specific question, enabling effective training of the frame selector.
@@
Without fine-tuning the downstream M-LLM for question answering, our frame selector enhances the video QA performance by reducing the noise introduced by irrelevant frames and sharpens the model’s focus on pertinent video segments. Our proposed approach shows significant improvements with various popular M-LLMs in video question answering tasks across short, medium (ActivityNet, NExt-QA) and long (EgoSchema) context video QA benchmarks. To summarize, our contributions are three folds:
@@
  • We propose spatial and temporal pseudo-labeling to generate importance scores for video frame select training.
@@ 2 Related Work
Multi-Modal Large Language Model (M-LLM) As Large Language Model (LLM) continue to demonstrate impressive abilities in language comprehension and reasoning [1, 36, 3, 39], interest is growing within the computer vision community to explore their potential for handling multi-modal inputs. Flamingo [2] demonstrates the capacity to process image and text inputs for a wide range of multi-modal tasks. BLIP-2 [13] introduces a Q-Former to map learned image features into the text embedding space of LLMs, while LLaVA [20] employs simple Multi Layer Perceptron (MLP) projector to align visual and textual features. Further research has focused on best practices for M-LLM, including areas such as dynamic high-resolution [21, 6], instruction-tuning data [15, 23], and different visual encoders [38, 44]. MM1 [28] and Idefics2 [11] provide comprehensive ablation studies on the design space of M-LLM.
@@
Video Frame Selection Before the rise of video M-LLM, language-aware video key-frame selection and localization had already attracted great interest [8, 22, 43, 7]. [5] optimized an end-to-end pipeline that uses ground truth question-answering labels to select a single key frame for downstream tasks. Lu et al. [24] uses a CLIP-like paradigm to train an additional transformer for visual and text feature alignment. Qian et al. [31] trains a video clip proposal model and the downstream QA model in a iterative manner respectively. Kim et al. [10] employs a semi-parametric retriever to obtain the key frames by comparing similarities between frame and language features.
Video M-LLM Frame Selection Several works propose to select key frames to improve video QA performance. SeViLA [50] prompts an M-LLM to obtain a relevance score to each frame, and use these scores to localize important frames. MVU [32] adopts a similar approach as SeViLA. However, a critical limitation of this kind of methods is the absence of temporal reasoning. Each frame is asserted independently without contextual information from other frames. Furthermore, this method produce expensive costs during inference. Every frame are required to estimated the importance score via a M-LLM. The longer the video the higher the cost. In contrast, our method outputs importance scores for all frames in a single pass with the compressed visual tokens, which reduces the computational costs and enables temporal reasoning. Koala [35] uses sparsely sampled key frames as conditions for processing subsequent visual tokens. ViLA [42] trains an end-to-end frame selection module to mask input frames. However, both of these methods lack text awareness: during inference, Koala’s subsequent visual token processing and ViLA’s frame masking do not account for the specific question posed about the video.
@@ 3 Method
This section introduces our video frame selector designed for efficient video-LLM QA. Section 3.1 discusses the motivation behind frame selection. Section 3.2 outlines the design details of the frame selector. Section 3.3 explains the generation of pseudo labels for training the frame selector. Section 3.4 describes the training process of the frame selector.
@@ 3.1 Rethinking Uniform Sampling in Video LLMs
A typical framework for video LLM   An n-frame framework is widely adopted in existing research on video M-LLM [14, 48, 54]. The number of frames in the input video, denoted by $T$, is variable. For example, a 3-minute video at 30 frames per second (FPS) contains $T=5400$ frames. The n-frame framework uniformly samples a fixed number of frames, $[x_{1},x_{2},\cdots,x_{n}]$, from the total $T$ frames, where each frame $x_{i}\in\mathbb{R}^{H\times W\times 3}$, with $H\times W$ representing the frame resolution, and typically $n\ll T$. A pre-trained visual encoder $f_{v}$ extracts visual features from n frames. These features are subsequently projected into the LLM’s space using an alignment projector $g_{a}$ and then flattened. Spatial pooling may also be applied to reduce the number of output tokens:
@@
Uniform sampling is not optimal  In the n-frame framework, as shown in Figure 2 (a), the input video is represented by $n\times m$ tokens where $m$ is the number of visual tokens per frame. For example, in LLaVA-NeXT-Video [54], where $n=32$ and $m=12\times 12$, this results in 4608 tokens. To reduce the computational cost of LLM inference, previous work has either chooses to reduce $n$ with sliding Q-Former [14] or $m$ [48, 17] by spatial pooling. However, both of them ignore the importance of reducing the number of key frames. In other words, to adaptively reduce $n$ before encoding. $n$ could be large especially in long videos. Denser uniform sampling makes a larger $n$, thereby reducing the efficiency of the video M-LLM. On the other hand, sampling fewer frames risks omitting crucial information. For example, sampling 32 frames from a 3-minute video means taking one frame every 6 seconds, potentially missing actions that occur within shorter time windows. In fact, most questions about a video can be answered using only a limited number of key frames. Inspired by this factor, we argue that adaptive frame selection according to question is more efficient and effective than uniform frame sampling.
@@ 3.2 Design of the Frame Selector
An ideal frame selector should be able to understand complex questions and analyze temporal relationships from the video input. To achieve this, we fine-tune an LLM to function as the frame selector. The frame selector utilizes the base LLM’s strong language comprehension and reasoning abilities to identify key frames in a video.
@@
$s=\textit{MLP}(e^{q}),s\in\mathbb{R}^{n}$ (5)
Figure 2 (b) presents an overview of the proposed architecture for the M-LLM-based frame selector. Instead of generating tokens that represents the frame selection, we append a learnable query vector at the end of the input sequence and supervisedly learn this query token. The hidden vector of this query token, $e^{q}$, is then served as the input to generate the $n$-dimensional importance vector $s$.
@@
2:Initialize the neighbor gap $\delta\leftarrow\text{int}(\nicefrac{{n}}{{4k}})$.
@@
Although dense uniform sampling is inefficient for video M-LLM, we keep using it at the beginning to maximally preserve the video information. To overcome the drawback of dense uniform sampling, we apply spatial pooling to reduce the token count per video frame before feeding the frames into the selector. In particular, the spatial pooling reduce the number of visual token to a smaller value, e.g., $m=3\times 3$ (9 tokens), which is substantially fewer than the tokens per frame used in existing video LLMs, e.g., $m=12\times 12$ (144 tokens). This design follows such an assumption: for video question answering, the model requires a substantial number of tokens per frame to capture visual details, but far fewer tokens are needed to determine whether a frame is important. A rough outline of the frame is sufficient. Our empirical results Table 6 also justify this assumption.
@@ 3.3 Pseudo Labels for the Frame Selector
Temporal Pseudo Labels A significant limitation of single-frame evaluation is the lack of temporal reasoning. For example, considering the question: “What did the man do after picking up his hat?”, the video content following the action of picking up the hat is crucial for answering this question. However, when generating the spatial pseudo labels, it only consider one frame without taking care of the temporal context. Therefore, the spatial labels don’t know what occur after the action.
@@
We find it challenging to generating floating-point scores for an extensive list of frames for an LLM. Consequently, we ask the model to produce a list of the index of most helpful frames (see appendix for the detailed prompt). Frames included in this list are assigned a score of 1, while those excluded receive a score of 0. Figure 3 (b) shows the pipeline.
@@ 3.4 Training of the Frame Selector
We consider a two-stage instruction-tuning procedure. In stage 1, we freeze the pre-trained vision and LLM backbones and train the parameters of the alignment projector $g_{a}$, the learned score query $X_{\text{score}}$ and the score projector $g_{s}$. Figure 2 (b) shows the trainable modules in the red boxes and the frozen modules in the blue boxes. Stage 1 training is optimized over the two tasks alternatively. We use the following two tasks:
  • Visual instruction tuning Recall $r$ in Equation 2 is the generated response from the LLM, the objective is the cross entropy loss between $r$ and the ground truth response. This tasks train the projector $g_{a}$ to aligns the visual features with the pre-trained LLM embedding space.
  • Importance score prediction Recall $s$ in Equation 5 is the output importance score for $n$ frames, the objective is the binary cross entropy loss between $s$ and the pseudo labels generated from Section 3.3. This task provide a good initialization for the score query and the score projector.
@@ 4 Experiments
We begin by outlining our experimental setup in Section 4.1. Then we demonstrate that our frame selector improves the performance of well-trained M-LLMs without changing their parameters by selecting better frames in Section 4.2. We also conduct ablation studies to demonstrate the effectiveness and efficiency of our frame selection framework and results are in Section 4.3. At last, we showcase some qualitative examples of frames selected from the video according to the question in Section 4.4.

### v2

@@ Abstract
Recent advances in Multi-Modal Large Language Models (M-LLMs) show promising results in video reasoning. Popular Multi-Modal Large Language Model (M-LLM) frameworks usually apply naive uniform sampling to reduce the number of video frames that are fed into an M-LLM, particularly for long context videos. However, it could lose crucial context in certain periods of a video, so that the downstream M-LLM may not have sufficient visual information to answer a question. To attack this pain point, we propose a light-weight M-LLM-based frame selection method that adaptively select frames that are more relevant to users’ queries. In order to train the proposed frame selector, we introduce two supervision signals (i) Spatial signal, where single frame importance score by prompting an M-LLM; (ii) Temporal signal, in which multiple frames selection by prompting Large Language Model (LLM) using the captions of all frame candidates. The selected frames are then digested by a frozen downstream video M-LLM for visual reasoning and question answering. Empirical results show that the proposed M-LLM video frame selector improves the performances various downstream video Large Language Model (video-LLM) across medium (ActivityNet, NExT-QA) and long (EgoSchema, LongVideoBench) context video question answering benchmarks.
@@ 1 Introduction
Figure 1: An example of our video frame selection for video QA. Compared to uniform sampling, ours has higher hit rate.
@@
Large Language Model (LLM) has revolutionized numerous domains in Artificial Intelligence (AI[34, 3, 41, 44]. In the past year, Multi-Modal Large Language Model (M-LLM) has significantly improved the performances of Vision-Language Models (VLM) to an unprecedented level in tasks such as image captioning and Visual Question Answering (VQA[2, 16, 4, 46]. To extend VQA to the temporal domain, video QA task requires a model to understand consecutive images along the time to answer a question, raising new challenges to the current M-LLM. One of the most critical challenges is the length of the video context, where a model needs to comprehend all frames in a video.
@@
To balance the trade-off between the capability of understanding all video frames and the available context length in a specific M-LLM, conventional practices [55, 61] rely on uniform sampling of frames. Frames are extracted at pre-defined intervals regardless of their relevance to the specific questions. Although such methods maximize the coverage of a video in time-axis, it introduces insufficient visual information. The uniformly sampled frames may be irrelevant or redundant, meanwhile some important frames are ignored. This fact prevents the model from focusing on key events in the video and sometimes increases the computational costs. Generally in video QA [59, 26], some specific frames are more likely to contain information relevant to the question. The conventional one-size-fits-all approach limits both the performance and practicality of M-LLMs, especially when working with long videos or resource-limited environments.
@@
To address these limitations, a straightforward idea is to select frames instead of uniform sampling [57, 35]. Focusing on the frames that are most helpful for the question, we can use significantly less number of frames without sacrificing the quality of video understanding. Illustrated as an example in Figure 1, frames that contains the most informative visuals can help downstream model to answer the question. To implement the above idea, we propose a light weight frame selector that employs fine-tuned version of an LLM. It makes use of M-LLM’s capability of multi-modality understanding to effectively capture the relevance between video frames and the question. We introduce two design choices to make the framework lightweight: (i) small LLMs are capable to understand complicated user questions; (ii) compress per video frame tokens to balance the long context trade-off.
@@
Although video QA tasks often require a large number of tokens per frame to capture content details, we hypothesize that determining frame importance does not require excessive tokens. Instead of leveraging all visual tokens of the frames, we apply an aggressive pooling-based token reduction on each video frame, which significantly improves the computational efficiency and increases the number of frames that the lightweight LLM selector can ingest.
@@
We also propose a training method for our LLM-based video frame selector. Since well-maintained video frame selection datasets for video QA are scarce, supervised training alone is not feasible. To address this issue, we adopt two pseudo-labeling strategies to estimate frame importance. The first strategy focuses on spatial understanding: we prompt a well-trained M-LLM using a video QA question, which then assigns importance scores to each frame. However, due to the limited context length of a M-LLM, it cannot evaluate all frames together from a temporal viewpoint. Therefore, our second strategy leverages an LLM to identify the top-k relevant frames using their captions. Instead of visual tokens, the LLM can analyze more frames simultaneously with caption. We integrate these two approaches to generate pseudo-labels that reflect frame importance relative to a specific question for effective training of the frame selector.
@@
Our frame-selector employs a plug-and-play design that requires no additional fine-tuning of the downstream M-LLM. It reduces noise caused by irrelevant frames, allowing the video-LLM to focus more effectively on the relevant content. Additionally, the selector only needs to be trained once and can subsequently enhance the video QA performance of multiple M-LLMs. We demonstrate significant improvements with several popular M-LLMs on video question-answering tasks across various benchmarks, including short-to-medium context (ActivityNet, NExt-QA) and long context (EgoSchema, VideoMME) scenarios. In summary, our contributions are threefold:
@@
  • We propose spatial and temporal pseudo-labeling to generate importance scores for video frame selector training.
@@ 2 Related Work
Multi-Modal Large Language Model (M-LLM) As Large Language Models (LLMs) continue to demonstrate impressive abilities in language comprehension and reasoning [1, 41, 3, 44], interest is growing within the computer vision community to explore their potential for handling multi-modal inputs. Flamingo [2] demonstrates the capacity to process image and text inputs for a wide range of multi-modal tasks. BLIP-2 [17] introduces a Q-Former to map learned image features into the text embedding space of LLMs, while LLaVA [25] employs simple Multi Layer Perceptron (MLP) projector to align visual and textual features. Further research has focused on best practices for M-LLM, including areas such as dynamic high-resolution [26, 6], instruction-tuning data [19, 28], and different visual encoders [43, 51]. MM1 [33] and Idefics2 [15] provide comprehensive ablation studies on the design space of M-LLM.
@@
Video Frame Selection Before the rise of video M-LLM, language-aware video key-frame selection and localization had attracted great interest [10, 27, 49, 7]. Buch et al. [5] optimized an end2end pipeline that uses ground truth question-answering labels to select a single key frame for downstream tasks. Lu et al. [29] uses a CLIP-like paradigm to train a transformer for visual and text feature alignment. Qian et al. [36] trains a video clip proposal model and the downstream QA model in a iterative manner. Kim et al. [14] employs a semi-parametric retriever to obtain the key frames by comparing similarities between frame and language features.
Video M-LLM Frame Selection Several works propose to select key frames to improve video QA performance. SeViLA [57] prompts an M-LLM to obtain a relevance score to each frame, and uses these scores to localize important frames. MVU [37] and adopts a similar approach as SeViLA. However, a critical limitation of this kind of methods [12, 22] is the absence of temporal reasoning. Each frame is assessed independently without contextual information from other frames. Furthermore, this method is expensive during inference. Importance score of every frame is estimated via a M-LLM. The longer the video the higher the cost. In contrast, our method outputs importance scores for all frames in a single pass with the compressed visual tokens, which reduces the computational costs and enables temporal reasoning. Koala [40] uses sparsely sampled key frames as conditions for processing subsequent visual tokens. ViLA [47] trains an end-to-end frame selection module to mask input frames. However, both of these methods lack text awareness: during inference, Koala’s subsequent visual token processing and ViLA’s frame masking do not account for the specific question posed about the video. Recent work such [50, 48] are not end-to-end methods.  Han et al. [11] finds frame selection is helpful for video chain-of-thought reasoning.
@@ 3 Method
This section introduces our video frame selector designed for efficient video-LLM QA. Section 3.1 disccuss our motivation. Section 3.2 outlines the design details of the frame selector. Section 3.3 explains the generation of pseudo labels for training the frame selector. Section 3.4 describes the training process of the frame selector.
@@ 3.1 Rethinking Uniform Sampling in Video LLMs
A typical framework for video LLM   An n-frame framework is widely adopted in existing research in video M-LLM [18, 55, 61]. The number of frames in the input video, denoted by $T$, is variable. For example, a 3-minute video at 30 frames per second (FPS) contains $T=5400$ frames. The n-frame framework uniformly samples a fixed number of frames, $[x_{1},x_{2},\cdots,x_{n}]$, from the total $T$ frames, where each frame $x_{i}\in\mathbb{R}^{H\times W\times 3}$, with $H\times W$ representing the frame resolution, and typically $n\ll T$. A pre-trained visual encoder $f_{v}$ extracts visual features from n frames. These features are subsequently projected into the LLM’s space using an alignment projector $g_{a}$ and then flattened. Spatial pooling may also be applied to reduce the number of output tokens:
@@
Uniform sampling is not optimal  In the n-frame framework, as shown in Figure 2 (a), the input video is represented by $n\times m$ tokens where $m$ is the number of visual tokens per frame. For example, in LLaVA-NeXT-Video [61], where $n=32$ and $m=12\times 12$, this results in 4608 tokens. To reduce the computational cost of LLM inference, previous work has either chosen to reduce $n$ with sliding Q-Former [18] or $m$ [55, 21] by spatial pooling, in other words, reducing $m$. However, both of them ignore the importance of reducing $n$, the number of intake frames, before encoding and $n$ could be large especially in long videos. Denser uniform sampling makes a larger $n$, thereby reducing the efficiency of the video M-LLM. On the other hand, sampling fewer frames risks omitting crucial information. For example, sampling 32 frames from a 3-minute video means taking one frame every 6 seconds, potentially missing actions that occur within shorter time windows. In fact, most questions about a video can be answered using only a limited number of key frames. Inspired by this factor, we argue that adaptive frame selection according to question is more efficient and effective than uniform frame sampling.
@@ 3.2 Design of the Frame Selector
An ideal frame selector should be able to understand complex questions and analyze temporal relevance in the video input. To achieve this, we fine-tune an LLM to function as the frame selector. The frame selector utilizes the base LLM’s strong language comprehension and reasoning abilities to identify key frames in a video.
@@
$s=\textit{MLP}(e^{q}),s\in\mathbb{R}^{n}\vspace{-12pt}$ (5)
Figure 2 (b) presents an overview of the proposed architecture for the M-LLM-based frame selector. Instead of generating tokens that represent the frame selection, we append a learnable query vector at the end of the input sequence and supervisedly learn this query token. The hidden vector of this query token, $e^{q}$, then serves as the input to generate the $n$-dimensional importance vector $s$.
@@
2:Initialize the neighbor gap $\delta\leftarrow\text{integer}(\nicefrac{{n}}{{4k}})$.
@@
Although dense uniform sampling is inefficient for video M-LLM, we keep using it at the beginning to maximally preserve the video information. To overcome the drawback of dense uniform sampling, we apply spatial pooling to reduces the token count per video frame before feeding the frames into the selector. In particular, the spatial pooling reduce the number of visual token to a smaller value, e.g., $m=3\times 3$ (9 tokens), which is substantially fewer than the tokens per frame used in existing video LLMs, e.g., $m=12\times 12$ (144 tokens). This design follows such an assumption: for video question answering, the model requires a substantial number of tokens per frame to capture visual details, but far fewer tokens are needed to determine whether a frame is important. A rough outline of the frame is sufficient. Our empirical results Table 7 also justify this assumption.
@@ 3.3 Pseudo Labels for the Frame Selector
Temporal Pseudo Labels A significant limitation of single-frame evaluation is the lack of temporal reasoning. For example, considering the question: “What did the man do after picking up his hat?”, the video content following the action of picking up the hat is crucial for answering this question. However, when generating the spatial pseudo labels, it only considers one frame without taking care of the temporal context. Therefore, the spatial labels don’t know what occurs after the action.
@@
We find it challenging to generate floating-point scores for an extensive list of frames for an LLM. Consequently, we ask the model to produce a list of the index of most helpful frames (see appendix Section A for details). Frames included in this list are assigned a score of 1, while those excluded receive a score of 0. Figure 3 (b) shows the pipeline.
@@ 3.4 Training of the Frame Selector
We consider a two-stage training procedure. In stage 1, we freeze the pre-trained vision and LLM backbones and train the parameters of the alignment projector $g_{a}$, the learned score query $X_{\text{score}}$ and the score projector $g_{s}$. Figure 2 (b) shows the trainable modules in the red boxes and the frozen modules in the blue boxes. Stage 1 training is optimized over the two tasks alternatively. We use below two tasks:
  • Visual instruction following Recall $r$ in Equation 2 is the generated response from the LLM, the objective is the cross entropy loss between $r$ and the ground truth response. This tasks trains the projector $g_{a}$ to aligns the visual features with the pre-trained LLM embedding space.
  • Importance score prediction Recall $s$ in Equation 5 is the importance score for $n$ frames, the objective is the binary cross entropy loss between $s$ and the pseudo labels generated from Section 3.3. This task provides a good initialization for the score query and the score projector.
@@ 4 Experiments
We begin by outlining our experimental setup in Section 4.1. Then we demonstrate that our frame selector improves the performance of well-trained M-LLMs without changing their parameters by selecting better frames in Section 4.2. We also conduct ablation studies to demonstrate the effectiveness and efficiency of our frame selection framework and results are in Section 4.3. At the end, we showcase some qualitative examples of frames selected from the video according to the question in Section 4.4.